Circular No. 7006

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998dr IN ANONYMOUS GALAXY J. Maza, Department of Astronomy, University of Chile, reports the

discovery of a supernova (B ∼ 18.0) by Roberto Antezana on a T-Max 400 film taken by L. E. Gonzalez with the Maksutov telescope at Cerro El Roble on Aug. 28.270 UT. SN 1998dr is located at α = 0^{h}09^{m}30^{s}.68, δ = −48^{o}50′39′′.1 (equinox 2000.0), which is 0′′.7 north of the nucleus of the host spiral galaxy. The new star is not present on a film obtained on July 31.310 (B > 20). SN 1998dr was confirmed by Maza at Las Campanas Observatory using the 2.5-m du Pont telescope on Sept. 2; a preliminary analysis of the spectrum shows this to be a type-Ia supernova, relatively near maximum.

SUPERNOVA 1998ds IN ANONYMOUS GALAXY Maza also reports the discovery of another supernova (B ∼ 18.5) by

Antezana on a film taken (as above) on Aug. 28.306 UT. SN 1998ds is located at α = 3^{h}17^{m}19^{s}.21, δ = −53^{o}10′58′′.0 (equinox 2000.0), which is 9′′.9 west and 2′′.9 north of the nucleus of the host spiral galaxy. The object is not present on a film obtained on July 31.382 (B > 20). SN 1998ds was confirmed by Maza with the du Pont 2.5-m telescope on Sept. 2; a preliminary analysis of the spectrum shows this to be a type-Ia supernova, relatively near maximum.

SUPERNOVA 1998dt IN NGC 945 T. Shefler, E. Halderson, M. Modjaz, J. Y. King, W. D. Li, R. R.

Treffers, and A. V. Filippenko, University of California at Berkeley, report their discovery of an apparent supernova during the course of the Lick Observatory Supernova Search (cf. IAUC 6627) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 1998dt was found and confirmed on unfiltered images taken on Sept. 1.4 (mag ∼ 17.7) and 2.5 UT (mag ∼ 17.3), respectively; it is located at α = 2^{h}28^{m}35^{s}.72, δ = −10^{o}32′59′′.8 (equinox 2000.0), which is ∼ 23′′.1 west and 39′′.5 south of the nucleus of NGC 945. A KAIT image obtained on Aug. 28.4 already shows a hint of the supernova, but an image taken on Aug. 23.4 shows no star at the position of SN 1998dt to a limiting mag of ∼ 19.0.

1998 September 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green